OpenStudy (anonymous):

evaluate the integral from 0 to 2pi, cos^(2)(6theta) dtheta

OpenStudy (anonymous):

You need to use the half angle formula which says: \[\cos^2(x)=\frac{1}{2}(1+\cos(2x))\] Applying this we have: \[\frac{1}{2}\int\limits_0^{2\pi}1+\cos(12 \theta)d \theta=\frac{1}{2}[\theta+\frac{1}{12}\sin(12\theta)]_0^{2\pi}\] You do the arithmetic :P

OpenStudy (anonymous):

omgsh thank u!

OpenStudy (anonymous):

No problem :P